OpenStudy (anonymous):

Q:The horizontal distance traveled by a projectile fire with initial angle theta and initial velocity 60 m/s can be modeled by the equation h=(7200/9.8) sin theta cos theta Choose the angle where the horizontal distance is at a maximum?" A: pi/6 pi/4 pi/2 pi/3" So is the equation have to be 60x(7200/9.8)sin theta cos theta? or graph it?

OpenStudy (anonymous):

can u help me here? what is the center of the circle that has a diameter whose endpoints are (2,7) and (-6, -1)? ans: (4,3), (-2, 4), or (-2,3)?

OpenStudy (anonymous):

4,4,3 I believe

OpenStudy (anonymous):

because its just looking for mid point of the diameter

OpenStudy (anonymous):

for your question or mine?

OpenStudy (anonymous):

your's its -2,3

OpenStudy (cwrw238):

the center is (2-6) / 2 , (7-1)/ 2

OpenStudy (cwrw238):

ladrybro is correct

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

find the distance between the points (-4,-5) and (3,-1)

OpenStudy (cwrw238):

the anggle givingthe maximum horizontal distance in ladybro's question is the value of theta where sin theta cos theta is a maximum

OpenStudy (cwrw238):

you can use calculus to find this value which is pi/4 ( 45 degrees)

OpenStudy (anonymous):

okay so input it into the theta the 45 degree?

OpenStudy (cwrw238):

to find the value of the max distance plug 45 degrees into your formula

OpenStudy (anonymous):

I got 367.3469.. with the equation above

OpenStudy (cwrw238):

h=(7200/9.8) sin theta cos theta is the correct formula not 60 * h=(7200/9.8) sin theta cos theta

OpenStudy (cwrw238):

h = (7200/9.8) * (1/sqrt2) * (1/sqrt2) = 3600 / 9.8

OpenStudy (cwrw238):

your answer is correct